OpenStudy (anonymous):

MEDAL AND FAN :D P(X>85)=P(X−62/10>85−62/10)=?

OpenStudy (xapproachesinfinity):

discrete probab?

OpenStudy (anonymous):

the full question is: In the city of Longskank, the mean weight is 62 and the standard deviation is 10. What is the probability that a person chosen at random weighs more than 85 kg? but i thought that that equation was how u solve it - i just didnt know how to work with it

OpenStudy (xapproachesinfinity):

why did you set this P(X−62/10>85−62/10)

OpenStudy (anonymous):

so do u know the answer @xapproachesinfinity???

OpenStudy (anonymous):

cuz i thought that that was easier

OpenStudy (xapproachesinfinity):

Not really, i forgot how to deal with this kind of probab just trying if i can remember something lol sorry

OpenStudy (anonymous):

sorry my internets being slow i didnt realize

OpenStudy (anonymous):

okayyy...

OpenStudy (xapproachesinfinity):

@ikram002p

OpenStudy (anonymous):

@kirbykirby maybe u could help me on this one too sorry

OpenStudy (kirbykirby):

Yes you did approach this correctly. You have \[P(X>85)=P\left(\frac{X-62}{10} >\frac{85-62}{10}\right)=P(Z>2.3) \] So now you can use a standard normal table to find this quantity, or use a calculator/program that has a function to find this

OpenStudy (anonymous):

do u know what the button would look like on the calc? I have a Ti-nspire here so that should have the function?

OpenStudy (kirbykirby):

um, I don't have a calculator that has these functions :( so I'm not sure which button you use

OpenStudy (anonymous):

ok thats cool

OpenStudy (kirbykirby):

did you ever use normal tables?

OpenStudy (anonymous):

not rly but im sure google would tell me?

OpenStudy (kirbykirby):

maybe if you search your calculator. What did you use before to find these probabilities? Otherwise, you can 1) learn to use normal tables 2) Try this here: http://www.wolframalpha.com/input/?i=P%28Z%3E2.3%29+where+Z~N%280%2C1%29

OpenStudy (xapproachesinfinity):

@kirbykirby where can i review this

OpenStudy (kirbykirby):

Um well if you need to use your calculator, I suggest going on the texas instruments website and search for your calculator model. You should be able to find the instructions for it as a PDF file. If you want to learn to use normal tables, you can try here: http://courses.wcupa.edu/rbove/Berenson/10th%20ed%20CD-ROM%20topics/section6_1.pdf

OpenStudy (anonymous):

thanks so much

OpenStudy (kirbykirby):

yw! :)